OpenStudy (anonymous):

second derivative of parametric equations: x=cos(t)**3 y=2sin(t)**2

OpenStudy (anonymous):

is it:\[x=cos(t)*3~; y=2sin(t)*2\]

OpenStudy (sirm3d):

\[\frac{ dy }{ dx }=\frac{ \frac{ dy }{ dt } }{ \frac{ dx }{ dt } }\] you need to compute each part of the fraction on the right side of the equation

OpenStudy (anonymous):

\[x=\cos ^{3}(t) ; y=2\sin ^{2}(t)\]

OpenStudy (anonymous):

second derivative of parametric equations: x=(cos^3(t)) and y=2(sin^2(t))

OpenStudy (sirm3d):

can you try finding \[\frac{ dy }{ dt }\] and \[\frac{ dx }{ dt }\]